Medical Research Corporation is expanding its research and production capacity to introduce a new line of products. Current plans call for the expenditure of $100 million on four projects of equal size ($25 million each), but different returns. Project A is in blood clotting proteins and has an expected return of 18 percent. Project B relates to a hepatitis vaccine and carries a potential return of 14 percent. Project C, dealing with a cardiovascular compound, is expected to earn 11.8 percent, and Project D, an investment in orthopedic implants, is expected to show a 10.9 percent return. The firm has $27,000,000 in retained earnings. After a capital structure with $27,000,000 million in retained earnings is reached (in which retained earnings represent 60 percent of the financing), all additional equity financing must come in the form of new common stock. Common stock is selling for $27 per share and underwriting costs are estimated at $2.8 if new shares are issued. Dividends for the next year will be $0.51 per share (D1), and earnings and dividends have grown consistently at 11% percent per year. The yield on comparative bonds has been hovering at 10 percent. The investment banker feels that the first $20,500,000 of bonds could be sold to yield 10 percent while additional debt might require a 2 percent premium and be sold to yield 12 percent. The corporate tax rate is 30 percent. Debt represents 40 percent of the capital structure. Based on the two sources of financing, what is the initial weighted average cost of capital? (Use Kd and Ke.) At what size capital structure will the firm run out of retained earnings? Round your response to the nearest whole dollar What will the marginal cost of capital be immediately after that point? Round your response to two decimal places At what size capital structure will there be a change in the cost of debt? Round your response to the nearest whole dollar What will the marginal cost of capital be immediately after that point? Round your response to two decimal places
Answer #1
Debt portion 40%
Equity portion 60%
Currently available finance through retained earning          27,000,000
Available debt financing without raising additional equity          45,000,000
=27mn/60%

Hence, if the firm wantes to raise capital beyond $45Mn, it will run out of retained earnings and need to raise new equity

Cost of debt
Bond yield 10%
First round of debt funding          20,500,000
This debt will correspond to 40% of total capital. Hence total capital available in debt+ equity
=20,500,000/40%          51,250,000
Hence, marginal WACC and cost of debt will change after capital=          51,250,000
Tax rate 30%
WACC for capital upto first change in WACC
Pre tax cost of debt 10%
Post tax cost of debt 7.00% =10%*(1-30%)
Cost of equity
(Gordon model)
g 11%
Cost of issue 2.8
Share price 27
net share price 24.2 =27-2.8
Dividend payout 0.51
Dividend yield 2.11% =0.51/24.2
P= G1/(r-g)
27=0.51/(r-11%)
r= 0.51/27+11%
r= 12.89%
Cost of equity 12.89%
WACC
=40%*7%+60%*12.89% 10.53%
WACC after the change in cost of debt
Cost of debt
Bond yield (above $20.5Mn) 12%
Pre tax cost of debt 12%
Post tax cost of debt 8.40% =12%*(1-30%)
Cost of equity 12.89%
WACC
=40%*8.40%+60%*12.89% 11.09%
Hence, WACC after change in cost of debt= 11.09%